### v1

@@ Abstract
Real-World Assets (RWAs) have recently attracted increasing attention as a means of bridging traditional financial instruments with decentralized infrastructures. By representing assets such as bonds, commodities, and real estate on blockchains, RWAs can enhance liquidity, broaden accessibility, and extend the scope of decentralized finance. Industry forecasts further suggest rapid growth of tokenized RWAs in the coming years, underscoring their potential role in the evolution of digital financial markets. However, when deployed across multiple blockchains, RWAs face challenges such as repeated authentication on different chains and inefficiency caused by multi-step settlement protocols. To address these issues, we present a cross-chain framework for RWAs that emphasizes identity management, authentication, and interaction. The framework integrates Decentralized Identifiers and Verifiable Credentials with customized attributes to support decentralized identification, and incorporates an authentication protocol based on Simplified Payment Verification to avoid redundant verification across chains. Furthermore, we design a cross-chain channel that enables the settlement of RWAs without requiring channel closure, thereby improving operational efficiency. We implement the framework and evaluate it through simulations, which confirm its feasibility and demonstrate improvements in efficiency for RWAs in cross-chain settings.
@@ 1 Introduction
Real-World Assets (RWAs) refer to the tokenization of off-chain assets such as government bonds, commodities, and real estate, enabling them to be represented and transacted on blockchain systems [5]. They are attracting increasing attention in Web3 finance: as of June 2025, the total market capitalization of global stablecoins is approximately \$250 billion, while the market size of RWA tokenization grew from \$8.6 billion at the beginning of 2025 to \$23 billion in the first half of the year, representing a 260% increase. Looking ahead, Citibank estimates that tokenized RWAs may reach a scale of around \$4 trillion by 2030, indicating their potential role in the future development of digital financial markets [7, 23].
Blockchain [28, 11, 9] constitutes the key infrastructure for the tokenization and circulation of RWAs, providing transparency and immutability for asset issuance and transfer. In recent years, blockchain systems have expanded rapidly. By June 2025, CoinMarketCap reported more than ten thousand active cryptocurrency platforms [6], suggesting that RWAs can be deployed across an increasing number of heterogeneous blockchains. Although such distribution facilitates broader adoption [17, 11, 22], the confinement of RWAs to single blockchains limits their financial utility, as assets cannot be seamlessly transferred or traded across different chains. This situation is analogous to traditional financial markets where banks are unable to conduct transactions across borders, thereby restricting the mobility and efficiency of capital. Fortunately, cross-chain technologies [2, 13, 12], widely regarded as a primary means of overcoming information silos between blockchains, may provide a feasible approach to enabling RWA interactions across heterogeneous systems. Nevertheless, existing research and practice on RWAs remain confined to single-chain contexts and have not systematically examined their cross-chain realization (shown in Figure 1). To address this limitation, we propose a cross-chain RWA framework that aims to resolve the main challenges inherent in cross-chain processes.
@@
Inefficiency in Cross-Chain Operations. Even within a single blockchain, operations involving RWAs already incur latency due to consensus. For instance, the two most prominent blockchain platforms, Bitcoin and Ethereum, support approximately 7 TPS and 15 TPS, which is significantly lower than traditional payment networks such as Visa, capable of processing several thousand TPS [19]. In cross-chain environments, this inefficiency is further amplified, since completing a cross-chain interaction typically requires multiple interdependent operations across heterogeneous ledgers. To ensure atomicity, decentralized protocols such as Hashed Timelock Contracts (HTLCs) are commonly employed; however, they generally involve four distinct on-chain operations, which substantially prolong completion time and increase operational costs. Cross-chain channels [13, 15] attempt to mitigate this limitation by integrating HTLC-based mechanisms with off-chain channels to improve efficiency. However, in existing designs, each settlement requires closing the channel and reopening a new one for subsequent interactions, which introduces additional costs and delays. Therefore, the central challenge is to design mechanisms that avoid repeated channel closure during settlement, thereby enabling more efficient cross-chain interoperability for RWAs.
To fill this gap, we propose a cross-chain framework for RWAs to redefine the workflow of RWAs, with particular emphasis on identification, cross-chain authentication, and cross-chain interaction. To ensure decentralization throughout the process, we adopt Decentralized Identifiers (DIDs) and Verifiable Credentials (VCs), which are widely used in Web3, and extend them with customized attributes tailored to represent RWAs. Building on Simplified Payment Verification (SPV), we design a cross-chain authentication protocol that prevents redundant verification of RWAs across multiple chains. Finally, we introduce a novel cross-chain channel scheme that enables the settlement of RWAs without requiring channel closure, thereby supporting more efficient cross-chain interactions.
@@
    To improve efficiency in cross-chain operations, we develop a novel channel scheme that supports the settlement of RWAs without requiring channel closure.
@@ 2 Related Works
In this section, we review related work on Real-World Assets (RWAs) from both industry and academia.
@@ 2.1 RWAs in Industry
RWAs have already been actively deployed across various industry sectors. MakerDAO [8] integrated tokenized U.S. Treasuries into its collateral framework, reinforcing the stability of decentralized finance. HSBC [1] introduced a gold-backed RWA token for retail investors in Hong Kong. Securitize [21] operates a regulated platform for issuing and trading RWAs, enabling compliant digital access to private companies and funds. Brickken [4] has tokenized over USD 250 million in RWAs across 14 countries. Goldman Sachs and BNY Mellon [3] partnered to tokenize money market fund shares on blockchain, aiming to improve accessibility and settlement efficiency.
@@ 2.2 RWAs in Academia
In academia, research on real-world assets has primarily focused on single-chain architectures and survey studies. Chen et al. [5] analyzed the security of RWAs, highlighting risks in KYC/AML processes, oracle designs, and collateral structures. The work in [16] investigated the connection between RWAs and escrow mechanisms and discussed potential directions for future development. Zhao et al. [29] proposed a scalable framework that integrates Ethereum staking, Layer-2 rollups, and DAO governance for RWA management. Xia et al. [26] studied tokenization approaches for RWAs with the aim of enhancing liquidity and improving asset management practices. LEGO [18] provided a quantitative methodology that maps historical failures in blockchain systems to current designs, including those involving RWAs.
@@ 2.3 Summary and Motivations
Recent developments indicate that RWAs have already been deployed across multiple sectors of the financial industry, with initiatives ranging from tokenized government bonds and money market funds to gold and real estate. At the same time, academic research has largely concentrated on single-chain architectures and survey studies. This contrast reveals a critical gap: although RWAs are increasingly deployed on different blockchain systems, they remain isolated within their respective environments, lacking mechanisms for interoperability or cross-chain recognition. Each implementation therefore operates in a silo, limiting the broader financial utility of tokenized assets. Motivated by this observation, we propose a cross-chain framework for RWAs that addresses the challenges of redundant authentication and low efficiency, and enables secure, decentralized interoperability across multiple blockchains.
@@ 3.1 Decentralized Identifiers
Decentralized Identifiers (DIDs) provide a decentralized identity mechanism for Web3, enabling users to prove ownership and control of their identities without relying on centralized authorities [24].
Definition 1 (DID Operations)
The process of a DID is represented by a tuple of polynomial-time algorithms $\text{DID}\overset{\text{def}}{=}(\mathsf{KeyGen},\mathsf{Create},\mathsf{Resolve},\mathsf{Update},\mathsf{Deactivate})$:
@@
  • $(did,doc)\leftarrow\mathsf{Create}(sk,pk)$. Given a key pair $(sk,pk)$, the algorithm outputs a new identifier diddid and its associated document docdoc.
@@
Decentralized Identifiers have been widely adopted in Web3 ecosystems and are formally standardized by the World Wide Web Consortium (W3C) [20].
@@ 3.2 Verifiable Credentials
Verifiable Credentials (VCs) provide issuer-signed statements about a DID subject that can be verified by cryptographic proofs without trusting the presenter [25].
Definition 2 (VC Operations)
@@
An important feature of VCs is selective disclosure, which allows a holder to reveal only specific attributes of a credential while keeping all other information hidden, thereby preserving privacy during verification [27].
@@ 3.3 Simple Payment Verification
SPV (Simple Payment Verification) enables a verifier to check whether a transaction is included in the blockchain without downloading full blocks [28].
Definition 3 (SPV)
@@ 3.4 Off-Chain Channel
Definition 4 (Off-Chain Channel)
@@ 3.5 Hashed Timelock Contracts
Hashed Timelock Contracts (HTLCs) are conditional payment contracts widely used in cross-chain atomic swaps and payment channel networks [13]. They ensure that either both parties complete the transaction successfully or both are refunded, preventing unilateral loss.
Definition 5 (HTLC)
HTLC can be represented by a tuple of polynomial-time algorithms $\text{HTLC}\overset{\text{def}}{=}(\mathsf{Lock},\mathsf{Unlock},\allowbreak\mathsf{Refund})$.
  • $\mathsf{Lock}$: The sender generates a random preimage $\rho$ and computes its hash $h(\rho)$. The asset is locked in a contract with condition $h(\rho)$ and a timeout $T$.
  • $\mathsf{Unlock}$: The receiver obtains $\rho$ before the timeout and submits it to the contract, which verifies $h(\rho)$ and transfers the locked asset to the receiver.
  • $\mathsf{Refund}$: If the receiver fails to provide $\rho$ before the timeout, the contract refunds the locked asset to the sender, ensuring no funds are lost.
HTLCs provide a cryptographic mechanism for atomic payments, ensuring fairness by either transferring assets upon correct disclosure of the preimage or refunding the sender in case of timeout.
In the following section, we leverage these techniques to design our proposed scheme.
@@
This section begins with an overview of xRWA, followed by a detailed description of the proposed scheme.
@@ 4.1 Overview
The process begins with the digital representation of off-chain assets such as real estate, stablecoins, and private credit. Traditional approaches rely on centralized custodians or special-purpose vehicles that create contractual mappings between physical assets and their digital forms. xRWA instead adopts Decentralized Identifiers (DIDs) and Verifiable Credentials (VCs) for decentralized identity management. By embedding customized attributes into VCs, the framework binds issuers and holders to asset records, ensuring legitimacy while reducing reliance on intermediaries. After tokenization, assets may need to be recognized across multiple blockchains. Existing methods often require repeated verification on each chain. xRWA introduces a cross-chain authentication protocol based on Simplified Payment Verification (SPV) proofs. By linking DIDs and VCs with tokenized assets, and applying SPV validation, authentication performed on one chain can be reused across others. Cross-chain transfers also require efficient settlement. xRWA proposes a cross-chain channel that supports settlement without channel closure. This allows frequent interactions to occur off-chain, with only final states recorded on-chain, reducing on-chain costs.
In the following, we provide a detailed description of each component of the framework, covering asset identification and tokenization, cross-chain authentication, and cross-chain interaction.
@@ 4.2 Identification and Tokenization
To realize the identification of RWAs, we construct $\mathsf{RWA\text{-}CC}$ (RWA Composite Credential) using decentralized identifiers (DIDs) and verifiable credentials (VCs). The credential follows a modular design with four components: $\mathsf{asset}$, $\mathsf{identity}$, $\mathsf{compliance}$, and $\mathsf{custody}$. Each component includes its own $\mathsf{sStatus}$ and $\mathsf{sProof}$ to support independent revocation and verification, while a top-level proof guarantees overall integrity.
@@
The $\mathsf{asset}$ component defines the core identification and tokenization of a real-world asset. The top-level $\mathsf{id}$specifies the identifier of the credential instance itself. Within the $\mathsf{asset}$ object, the field $\mathsf{assetId}$ assigns a DID-based unique identifier to the asset. The field $\mathsf{assetType}$ indicates the broad category of the real-world asset, while $\mathsf{category}$ provides a finer classification within that type, for example distinguishing $\mathsf{Residential}$ from $\mathsf{Commercial}$ property. The $\mathsf{classDid}$ links the asset to a class definition maintained by an authoritative registry, enabling semantic grouping and interoperability across assets of the same class. A crucial element is $\mathsf{tokenBinding}$, which establishes the binding between the off-chain asset record and its tokenized representation on a blockchain. It records the token standard, the target chain, the contract address, and the token identifier in a unified structure.
The $\mathsf{identity}$ component extends the description of an asset beyond its basic identifier. The field $\mathsf{schemaVersion}$ specifies the version of the identity schema, while $\mathsf{identitySchema}$ provides a reference URI that defines its structure. The $\mathsf{identifiers}$ array records registration information from official registries; each entry contains details such as the registration scheme, the identifier value, the jurisdiction, and the issuing authority. The $\mathsf{taxonomies}$ field captures standardized classification codes together with descriptive labels, ensuring interoperability across different cataloguing systems. Spatial information is anchored by $\mathsf{spatialFootprint}$, which encodes the asset’s geographic location in GeoJSON format, including its geometry and granularity [14]. Additional evidence is attached through $\mathsf{documents}$, which may include deeds, certificates, or licenses, each linked to cryptographic hashes and media metadata. The $\mathsf{relations}$ field links the asset to other entities or classes, while $\mathsf{attributes}$ provide structured factual details such as size or physical characteristics. The $\mathsf{custom}$ object further allows local extensions, for example policy tags or IoT-based identifiers. Finally, the component includes $\mathsf{sStatus}$ for revocation or suspension tracking and $\mathsf{sProof}$ for cryptographic verification.
The $\mathsf{compliance}$ component records regulatory information that governs the circulation of the asset. The field $\mathsf{licenseId}$ identifies the license or permit under which the asset is authorized, while $\mathsf{sellableRegions}$ lists the jurisdictions where transfer is legally allowed. Restrictions such as $\mathsf{NoCrossBorderSale}$ are specified in $\mathsf{restrictions}$, and the effective period is defined by $\mathsf{effectiveFrom}$ and $\mathsf{effectiveTo}$. The $\mathsf{regulatorDid}$ provides the decentralized identifier of the authority that endorses or enforces these conditions. To ensure trustworthy usage, the component also includes $\mathsf{sStatus}$ for revocation or suspension tracking and $\mathsf{sProof}$ for independent cryptographic verification.
The $\mathsf{custody}$ component captures information related to the safekeeping of the asset. It specifies the custodian through $\mathsf{custodianDid}$, together with the physical or organizational $\mathsf{location}$ where the asset is maintained. The $\mathsf{policy}$ field encodes applicable standards such as security or compliance frameworks, and $\mathsf{auditCycleDays}$ defines the frequency of regular audits. Insurance arrangements are linked through $\mathsf{insurancePolicyRef}$. Similar to the compliance component, $\mathsf{sStatus}$ is included to manage revocation or suspension, and $\mathsf{sProof}$ ensures verifiable integrity.
@@
After an RWA is identified through $\mathsf{RWA\text{-}CC}$, the next step is to tokenize it on a blockchain. The process builds on the algorithms for decentralized identifiers (DIDs) and verifiable credentials (VCs) introduced in Section 3. The issuer $\mathcal{I}$ generates its decentralized identity using $\mathsf{DID.KeyGen}$ and obtains a DID $did_{\mathcal{I}}$. The asset $a$ is also assigned a DID $did_{a}$, which is included in the $\mathsf{asset}$ component of $\mathsf{RWA\text{-}CC}$. The issuer then executes $\mathsf{VC.Issue}$ to generate a credential $\mathsf{cred}$ that encapsulates the attributes of $a$, including $\mathsf{identity}$, $\mathsf{compliance}$, and $\mathsf{custody}$. This binds the off-chain asset to its verifiable description under the issuer’s DID. To make the asset tradable, the credential is linked to an on-chain token via $\mathsf{tBinding}$, which specifies the token standard (e.g., ERC-721), the blockchain, the contract address, and the identifier. This ensures that smart contracts interacting with the token can invoke $\mathsf{VC.Verify}$ against the issuer’s DID and status list to check its validity. Finally, when a holder $\mathcal{U}$ presents the token in applications, only the necessary attributes are disclosed through $\mathsf{VC.Prove}$, ensuring selective transparency while maintaining compliance guarantees.
1"id": "did:ion:EiAa...credential123",
@@
3 "assetId": "did:ion:EiAa...asset123",
4 "assetType": "RealEstate",
5 "category": "Residential",
6 "classDid": "did:web:issuer.example.org:class:RE-RESIDENCE",
7 "tokenBinding": {
8 "standard": "ERC-721",
9 "chain": "eip155:1",
10 "contract": "0xBC4CA0EdA7647A8aB7C2061c2E118A18a936f13D",
11 "tokenId": "1234"
@@
15 "schemaVersion": 1,
16 "identitySchema": "https://example.org/schemas/rwa-identity-v2.json",
@@
19 "identifierScheme": //Identification System,
20 "identifierValue": "example-2024-000123",
21 "jurisdiction": //Jurisdiction Code,
22 "issuingAuthority": //Issuing or Registration Agency

### v3

@@ Abstract
Real-World Assets (RWAs) serve as a bridge between traditional financial instruments and decentralized infrastructures. By representing assets such as bonds, commodities, and real estate on blockchains, RWAs can extend the scope of decentralized finance. Industry forecasts further indicate rapid growth in tokenized RWAs after 2025, underscoring their potential role in the evolution of digital financial markets. However, in the current multi-chain environment, RWAs face challenges such as repeated authentication across multiple chains and inefficiencies arising from multi-step settlement protocols. To address these issues, we present a cross-chain framework for RWAs that emphasizes identity management, authentication, and cross-chain interaction. The framework integrates Decentralized Identifiers and Verifiable Credentials with customized attributes to support decentralized identification, and incorporates an authentication protocol based on Simplified Payment Verification to avoid redundant verification across chains. Furthermore, this paper adopts a cross-chain channel that supports efficient RWA settlements, and we refine its design so that the channel does not need to be closed immediately after each settlement, thereby reducing on-chain cost. We implement the framework and evaluate its performance via simulations, which confirm its feasibility and demonstrate improvements in efficiency for RWAs in cross-chain settings.
@@ 1 Introduction
Real-World Assets (RWAs) refer to the tokenization of off-chain assets such as government bonds, commodities, and real estate, enabling them to be represented and transacted on blockchain systems [5]. They are attracting increasing attention in Web3 finance. As of June 2025, the total market capitalization of global stablecoins is approximately \$250 billion, while the market size of RWA tokenization grew from \$8.6 billion at the beginning of 2025 to \$23 billion in the first half of the year, representing a 260% increase. Looking ahead, Citibank estimates that tokenized RWAs may reach a scale of around \$4 trillion by 2030, indicating their potential role in the future development of digital financial markets [7, 23].
Blockchain [28, 12, 10] constitutes the key infrastructure for the tokenization and circulation of RWAs, providing transparency and immutability for asset issuance and transfer. In recent years, blockchain systems have expanded rapidly. By June 2025, CoinMarketCap reported more than ten thousand active cryptocurrency platforms [6], suggesting that RWAs can be deployed across an increasing number of heterogeneous blockchains. Although such distribution facilitates broader adoption [17, 12, 22], the confinement of RWAs to single blockchains limits their financial utility, as assets cannot be traded across different chains. This situation is analogous to traditional financial markets where banks are unable to conduct transactions across borders, thereby restricting the mobility and efficiency of capital. Fortunately, cross-chain technologies [2, 14, 13], regarded as a primary means of overcoming information silos between blockchains, may provide a feasible approach to enabling RWA interactions across heterogeneous systems. Nevertheless, existing research and practice on RWAs remain confined to single-chain contexts and have not systematically examined their cross-chain realization (shown in Figure 1). To address this limitation, we propose a cross-chain RWA framework that aims to resolve the main challenges inherent in cross-chain processes.
@@
Inefficiency in Cross-Chain Operations. Even within a single blockchain, operations involving RWAs already incur latency due to consensus. For instance, the two most prominent blockchain platforms, Bitcoin and Ethereum, support approximately 7 TPS and 15 TPS, which is significantly lower than traditional payment networks such as Visa, that can process several thousand TPS [19]. In cross-chain environments, this inefficiency is further amplified, since completing a cross-chain interaction typically requires multiple interdependent operations across heterogeneous ledgers. To ensure atomicity, decentralized protocols such as Hashed Timelock Contracts (HTLCs) are commonly employed; however, they generally involve four distinct on-chain operations, which substantially prolong completion time and increase operational costs. Cross-chain channels [14, 15] attempt to mitigate this limitation by integrating HTLC-based mechanisms with off-chain channels to improve efficiency. However, in existing designs, each settlement requires closing the channel and opening a new one for subsequent interactions, which introduces additional costs and delays. Therefore, the central challenge is to adopt cross-chain channels that support efficient interoperability for RWAs. Moreover, it is necessary to design mechanisms that avoid repeated channel closures during settlement, thereby enabling lower on-chain costs.
To fill this gap, we propose a cross-chain framework for RWAs to redefine the workflow of RWAs, with particular emphasis on identification, cross-chain authentication, and cross-chain interaction. To ensure decentralization throughout the process, we adopt Decentralized Identifiers (DIDs) and Verifiable Credentials (VCs), which are widely used in Web3, and extend them with customized attributes tailored to represent RWAs. Building on Simplified Payment Verification (SPV), we design a cross-chain authentication protocol that prevents redundant verifications of RWAs across multiple chains. Finally, we introduce a cross-chain channel scheme that enables the settlement of RWAs without requiring the channel to be closed after each settlement, thereby supporting efficient cross-chain interactions with lower costs.
@@
    We employ a cross-chain channel to achieve more efficient RWA settlement. We further redesign the channel workflow to reduce operational overhead by eliminating the repeated opening and closing of channels.
@@ 2 Related Works
In this section, we review related work on RWAs from both industry and academia.
@@ 2.1 RWAs in Industry
RWAs have already been actively deployed in various industry sectors. MakerDAO [8] integrated tokenized U.S. Treasuries into its collateral framework, reinforcing the stability of decentralized finance. HSBC [1] introduced a gold-backed RWA token for retail investors in Hong Kong. Securitize [21] operates a regulated platform for issuing and trading RWAs, enabling compliant digital access to private companies. Brickken [4] has tokenized RWAs worth over USD 250 million across 14 countries. Goldman Sachs and BNY Mellon [3] partnered to tokenize shares of money market funds on blockchains, aiming to improve accessibility and settlement efficiency.
The RWA tokenization market has grown by more than 380% from 2022 to 2025 [9]. These deployments now span heterogeneous blockchain infrastructures such as Ethereum, Polygon, Solana, and permissioned networks like Canton, resulting in RWA assets being distributed across isolated blockchain environments without shared standards for interoperability or cross-chain recognition.
@@ 2.2 RWAs in Academia
In academia, research on real-world assets has primarily focused on single-chain architectures and survey studies. Chen et al. [5] analyzed the security of RWAs, highlighting risks in KYC/AML processes, oracle designs, and collateral structures. The work in [16] investigated the connection between RWAs and escrow mechanisms, and discussed potential directions for future development. Zhao et al. [29] proposed a scalable framework that integrates Ethereum staking, Layer-2 rollups, and DAO governance for RWA management. Xia et al. [26] studied tokenization approaches for RWAs with the aim of enhancing liquidity and improving asset management practices. LEGO [18] provided a quantitative methodology that maps historical failures in blockchain systems to current designs, including those involving RWAs.
In academia, current research on real-world assets assumes a single underlying blockchain and lacks mechanisms for interoperability across heterogeneous chains. Existing studies do not address how RWAs deployed on different blockchain platforms can operate in multi-chain environments, resulting in fragmented deployments across multiple chains.
@@ 2.3 Summary and Motivations
Recent developments indicate that RWAs have already been deployed across multiple sectors of the financial industry, with initiatives ranging from tokenized government bonds and money market funds to gold and real estate. Meanwhile, academic research has largely concentrated on single-chain architectures and survey studies. This contrast reveals a critical gap: although RWAs are deployed on different blockchain systems, they remain isolated within their respective environments. Each implementation therefore operates in a silo, limiting the broader financial utility of tokenized assets. Motivated by this observation, we propose a cross-chain framework for RWAs that addresses the challenges of redundant authentication and low efficiency, and enables secure interoperability across multiple blockchains.
@@ 3.1 Decentralized Identifiers
DIDs provide a decentralized identity mechanism for Web3, enabling users to prove ownership and control of their identities without relying on centralized authorities [24].
Definition 1 (DID Operations).
The operations of a DID are represented as a tuple of polynomial-time algorithms $\text{DID}\overset{\text{def}}{=}(\mathsf{KeyGen},\mathsf{Create},\mathsf{Resolve},\mathsf{Update},\mathsf{Deactivate})$:
@@
  • $(did,doc)\leftarrow\mathsf{Create}(pk,sk)$. Given a key pair $(pk,sk)$, the algorithm outputs a new identifier diddid and its associated document docdoc.
@@
DIDs have been widely adopted in Web3 ecosystems and are formally standardized by the World Wide Web Consortium (W3C) [20].
@@ 3.2 Verifiable Credentials
VCs provide issuer-signed statements about a DID subject that can be verified by cryptographic proofs without trusting the presenter [25].
Definition 2 (VC Operations).
@@
An important feature of VCs is selective disclosure, which allows a holder to reveal only specific attributes of a credential while keeping all other information hidden [27].
@@ 3.3 Simple Payment Verification
SPV enables a verifier to verify whether a transaction is included in the blockchain without downloading full blocks [28].
Definition 3 (SPV).
@@ 3.4 Off-Chain Channel
Definition 4 (Off-Chain Channel).
@@ 3.5 Hashed Timelock Contracts
HTLCs are conditional payment contracts widely used in cross-chain atomic swaps [14]. They ensure that either both parties complete the transaction successfully or both are refunded, preventing unilateral loss.
Definition 5 (HTLC).
HTLC consists of three polynomial-time algorithms $\mathsf{HTLC}\allowbreak\overset{\mathsf{def}}{=}(\mathsf{Lock},\mathsf{Unlock},\mathsf{Refund})$. It operates on two blockchains $\mathbb{B}_{0}$ and $\mathbb{B}_{1}$, where both parties deploy contracts with the same hash value and coordinated timeouts.
  • $\mathsf{Lock}$: The initiator $P_{0}$ samples a random preimage $\rho$ and computes its hash $h(\rho)$. $P_{0}$ locks an asset on $\mathbb{B}_{0}$ in a contract that specifies $h(\rho)$ and a timeout $T_{0}$. The counterparty $P_{1}$ then deploys a corresponding contract on $\mathbb{B}_{1}$ using the same hash value and a shorter timeout $T_{1}<T_{0}$.
  • $\mathsf{Unlock}$: To redeem the asset on $\mathbb{B}_{1}$, $P_{0}$ reveals $\rho$ to the contract on $\mathbb{B}_{1}$ before $T_{1}$. Since all blockchain data are public, $P_{1}$ learns $\rho$ from this on-chain revelation and submits it to the contract on $\mathbb{B}_{0}$ before $T_{0}$ to unlock the asset. Both contracts verify that the submitted preimage satisfies $h(\rho)$ before releasing funds.
  • $\mathsf{Refund}$: If either party does not obtain $\rho$ before the corresponding timeout, the contract on $\mathbb{B}_{0}$ (resp. $\mathbb{B}_{1}$) refunds the asset to $P_{0}$ (resp. $P_{1}$), ensuring no loss in the event of failed coordination.
HTLCs implement cross-chain payments by enforcing that assets are released only when the correct preimage is disclosed, and otherwise returned to the sender after the timeout, thereby ensuring atomicity.
In the following section, we adopt these techniques to design our proposed scheme.
@@
This section begins with an overview of xRWA with its design goals, followed by a detailed description of the proposed scheme.
@@ 4.1 Overview
The process begins with the digital representation of off-chain assets such as real estate, stablecoins, and private credit assets. Traditional approaches rely on centralized custodians or special-purpose vehicles that create contractual mappings between physical assets and their digital forms. xRWA adopts DIDs and VCs for decentralized identity management. By embedding customized attributes into VCs, the framework binds issuers and holders to asset records, ensuring legitimacy while reducing the reliance on intermediaries. After tokenization, assets may need to be recognized across multiple blockchains. Existing methods often require repeated verification on each chain. xRWA introduces a cross-chain authentication protocol based on SPV proofs. By linking DIDs and VCs with tokenized assets, and applying SPV validation, authentication performed on one chain can be reused across others. Cross-chain transfers also require efficient settlement. xRWA proposes a cross-chain channel that supports settlement without channel closures. This allows frequent interactions to occur off-chain, with only final states recorded on-chain, reducing on-chain costs.
4.2 Design Goals
The security objective of xRWA is determined by two core guarantees that capture its correctness in multi-chain environments.
Cross-chain authentication correctness. A blockchain participating in xRWA may authenticate an RWA either (i) directly on its own ledger or (ii) by accepting an authentication result originating from another chain. For the second case, correctness requires that a chain $C_{t}$ accept such a cross-chain authentication claim only if it corresponds to a genuine and confirmed authentication record on the source chain $C_{s}$ according to the protocol rules. In particular, no PPT adversary should be able to fabricate cross-chain evidence so as to make $C_{t}$ accept an RWA based on a non-existent or invalid authentication record.
Let $\mathsf{AuthAcc}_{C_{t}}(a)$ denote the event that chain $C_{t}$ accepts asset $a$ via a cross-chain authentication claim from $C_{s}$, and let $\mathsf{AuthRec}_{C_{s}}(a)$ denote the existence of a valid and confirmed authentication record for $a$ on $C_{s}$. Authentication correctness requires that for every PPT adversary $\mathcal{A}$,
$\Pr\!\left[\mathsf{AuthAcc}_{C_{t}}(a)\;\wedge\;\neg\,\mathsf{AuthRec}_{C_{s}}(a)\right]\leq\mathsf{negl}(\lambda),$
meaning that a cross-chain claim should never be accepted unless it is consistent with an actual authentication on the source chain.
Cross-chain settlement atomicity. When two blockchains engage in an RWA settlement procedure, the resulting state changes must be all or nothing. If settlement succeeds, both ledgers must apply the corresponding asset and payment updates. If settlement fails, both ledgers must retain their original states, and no honest party should end up with a partially executed outcome.
Let $\mathsf{Com}_{1}$ and $\mathsf{Com}_{2}$ denote the events that chains $C_{1}$ and $C_{2}$, respectively, apply the settlement updates for a given cross-chain interaction. Atomicity requires that for every PPT adversary $\mathcal{A}$,
$\Pr\!\left[\mathsf{Com}_{1}\;\oplus\;\mathsf{Com}_{2}\right]\leq\mathsf{negl}(\lambda),$
where $\oplus$ denotes exclusive-or. Thus, it is infeasible for an adversary to cause exactly one chain to commit while the other does not, or to produce any partial or inconsistent settlement outcome.
Definition 6 (RWA Security).
A cross-chain RWA protocol $\Pi$ is said to satisfy RWA security if it ensures both (i) cross-chain authentication correctness and (ii) cross-chain settlement atomicity. Namely, for every PPT adversary $\mathcal{A}$, the probability that $\mathcal{A}$ causes a chain to accept a fabricated cross-chain authentication claim is negligible, and the probability that $\mathcal{A}$ produces an inconsistent or partially executed cross-chain settlement across two blockchains is negligible in the security parameter $\lambda$.
@@ 4.3 Identification and Tokenization
To realize the identification of RWAs, we construct $\mathsf{RWA\text{-}CC}$ (RWA Composite Credential) using DIDs and VCs. The credential follows a modular design with four components: $\mathsf{asset}$, $\mathsf{identity}$, $\mathsf{compliance}$, and $\mathsf{custody}$. Each component includes its own $\mathsf{sStatus}$ and $\mathsf{sProof}$ to support independent revocation and verification, while a top-level proof guarantees overall integrity.
@@
The $\mathsf{asset}$ component defines the core identification and tokenization of a real-world asset. The top-level $\mathsf{id}$ specifies the identifier of the credential instance itself. Within the $\mathsf{asset}$ object, the field $\mathsf{assetId}$ assigns a DID-based unique identifier to the asset. The field $\mathsf{assetType}$ indicates the broad category of the real-world asset, while $\mathsf{category}$ provides a finer classification within that type, for example distinguishing $\mathsf{Residential}$ from $\mathsf{Commercial}$ property. The $\mathsf{classDid}$ links the asset to a class definition maintained by an authoritative registry, enabling semantic grouping and interoperability across assets of the same class. A crucial element is $\mathsf{tBinding}$, which establishes the binding between the off-chain asset record and its tokenized representation on a blockchain. It records the token standard, the target chain, the contract address, and the token identifier in a unified structure.
The $\mathsf{identity}$ component extends the description of an asset beyond its basic identifier. The field $\mathsf{schemaVersion}$ specifies the version of the identity schema, while $\mathsf{identitySchema}$ provides a reference URI that defines its structure. The $\mathsf{identifiers}$ array records registration information from official registries: each entry contains details such as the registration scheme, the identifier value, the jurisdiction, and the issuing authority. The following components may incorporate attributes specifically designed to uniquely identify various categories of RWAs, such as location, applicable standards, associated documentation, or other distinctive characteristics. These attributes record the intrinsic properties of the RWA and establish a one-to-one correspondence between the VC and the RWA entity.
The $\mathsf{compliance}$ component defines the regulatory framework governing an asset’s circulation, including its authorizing license, permitted jurisdictions, applicable restrictions, validity period, and the regulator’s identifier. Similarly, the $\mathsf{custody}$ component records safekeeping information: custodian identity, storage location, operational policy, audit frequency, and insurance reference.
@@
After an RWA is identified through $\mathsf{RWA\text{-}CC}$, the next step is to tokenize it on a blockchain. The process builds on the algorithms for DIDs and VCs introduced in Section 3. The issuer $\mathcal{I}$ generates its decentralized identity using $\mathsf{DID.KeyGen}$ and obtains a DID $did_{\mathcal{I}}$. The asset $a$ is also assigned a DID $did_{a}$, which is included in the $\mathsf{asset}$ component of $\mathsf{RWA\text{-}CC}$. The issuer then executes $\mathsf{VC.Issue}$ to generate a credential $\mathsf{cred}$ that encapsulates the attributes of $a$, including $\mathsf{identity}$, $\mathsf{compliance}$, and $\mathsf{custody}$. This binds the off-chain asset to its verifiable description under the issuer’s DID. To make the asset tradable, the credential is linked to an on-chain token via $\mathsf{tBinding}$, which specifies the token standard (e.g., ERC-721), the blockchain, the contract address, and the identifier. This ensures that smart contracts interacting with the token can invoke $\mathsf{VC.Verify}$ against the issuer’s DID and status list to check its validity. Finally, when a holder $\mathcal{U}$ presents the token in applications, only the necessary attributes are disclosed through $\mathsf{VC.Prove}$, ensuring selective transparency while maintaining compliance guarantees.
1"id": "<did:example:subject>",
@@
3 "assetId": "<did:example:asset>",
4 "assetType": "<string>"
5 "category": "<string>"
6 "classDid": "<did:example:class>",
7 "tBinding": {
8 "standard": "<string>(e.g.,’ERC-721’)",
9 "chain": "<string>(e.g.,’eip155:1’)",
10 "contract": "<string>",
11 "tokenId": "<string>"
@@
15 "schemaVersion": <number>,
16 "identitySchema": "<URL>",
@@
19 "identifierScheme": "<string>",
20 "identifierValue": "<string>",
21 "jurisdiction": "<string>",
22 "issuingAuthority": "<string>"